OpenStudy (butterflyhope):

URGENT *WILL MEDAL*: Line AB has endpoint A(5,−4) and midpoint M(1,2). Which ordered pair are the coordinates of B? Thanks c:

jimthompson5910 (jim_thompson5910):

Let B be the point (x,y)

jimthompson5910 (jim_thompson5910):

A = (5,-4) B = (x,y) the midpoint of AB is found by adding the corresponding coordinates and dividing by 2 x coordinate of midpoint: (5+x)/2 y coordinate of midpoint: (-4+y)/2

jimthompson5910 (jim_thompson5910):

so we have these two equations (5+x)/2 = 1 (-4+y)/2 = 2 I'll let you solve

OpenStudy (butterflyhope):

I still don't know how to solve :/ :(

OpenStudy (butterflyhope):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

does it help to multiply both sides of (5+x)/2 = 1 by 2 to get (5+x)/2 = 1 2*(5+x)/2 = 2*1 5+x = 2 do you see where to go from here?

OpenStudy (butterflyhope):

So the point is (-3, 8)? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

very good